# June 13, 2024

### v2

@@ 4.1 Experiment Setup
Table 4: The comparison of AE’s visual quality between DifAttack++ and MCGSquare.
@@ 4.2 The Close-set Scenario
Table 2 and 3 present the ASR and average query numbers (Avg.Q) when attacking victim models from the “simple group" or “complex group". For targeted attacks, we randomly select a target class (864) and 1,000 images correctly classified by the victim model, excluding those belonging to the selected class, as the test set. It is demonstrated that on average, our DifAttack++ surpasses all compared methods in both ASR and Avg.Q. Moreover, in both targeted and untargeted attacks, DifAttack++ usually achieves the same or higher ASR than the second-best method, MCGSquare, while using much fewer Avg.Q especially for the challenging targeted attacks and powerful victim models. For example, in the targeted attacks for the “simple group" (see Table 2), DifAttack++ improves the average ASR by an average of 1% while reducing query counts by 37.4%. For the complex group (see Table 3), DifAttack++ achieves a higher ASR in targeted attacks while reducing the Avg.Q on average by nearly 1,000, and in untargeted attacks, it further decreases the number of queries by 41.5%. Besides the overall attack efficiency improvement, DifAttack++ also significantly outperforms MCGSquare in terms of the visual quality of AEs, which makes DifAttack++ less perceptible and more practical. Now we present more details in the following subsection.
4.2.2 DifAttack++’s Advantages over MCGSquare.
@@
AE’s Qualitative Visual Quality. Due to the algorithm charactistics, SquareAttack results in severe visual distortions (e.g., colored blocks or stripes) in their generated AEs (see more examples in Andriushchenko et al (2020)). So AEs generated by MCGSquare, which enhances the preliminary SquareAttack method by using meta learning and historic query information, inevitably exhibits such severe visual perturbations, as can be seen from Fig. 2. In contrast, our DifAttack++ achieves efficient attacks while ensuring much subtle visual perturbation in the generated AEs, making them imperceptible to human eyes. This is because we effectively disentangle the visual and adversarial features of images, and ensure that the VFs of the image remain almost unchanged during the generation of AEs, while updating only the AFs. AE’s Quantitative Visual Quality. We also evaluate the average image quality metrics of a set of AEs generated by our DifAttack++ or MCGSquare, when comparing these AEs with their associated clean images. These AEs are randomly produced from the ImageNet test dataset and successfully fool the ResNet18 model. From the result summarized in Table 4 we can see that in terms of the four metrics considered, the quality of AEs’ generated by DifAttack++ significantly surpasses that of MCGsquare, e.g. over 2.5 higher PSNR values.
Figure 3: The comparison of Avg.Q in untargeted attacks on the ImageNet test dataset between DifAttack and DifAttack++ w/ $\mathbf{x}$, in which our DifAttack++ method begins the attack stage with clean images to maintain fairness.
4.2.3 DifAttack++’s Advantages over DifAttack.
From Table 2-3, it is demonstrated that our DifAttack++ outperforms DifAttack significantly in terms of ASR and query efficiency. Specifically, when attacking relatively simple victim models, as shown in Table 2, DifAttack++ boosts the Avg.Q of DifAttack by 56.7% and 26.2% in average in untargeted and targeted attacks, respectively. When the complexity of victim models increases (see Table 3), the gains of DifAttack++ are even more pronounced, with 68.4% and 35.9% enhancements in Avg.Q, meanwhile obtaining 13.8% higher ASR in targeted attacks. In this subsection, we illustrate that DifAttack++ also greatly enhances the training efficiency and test stability. In Fig. 3, we depict the changes of Avg.Q under untargeted attacks by DifAttack++ or DifAttack, where attacks are conducted on the same ImageNet test set and four different victim models over training time. It is worth mentioning that, for fair comparison, in this close-set scenario, we use the results of DifAttack++ with clean images as the initial image for the attack stage, denoted as DifAttack++ w/ $\mathbf{x}$, to align with DifAttack. As shown in Fig. 3(a) and (b), it is indicated that, generally, models trained with DifAttack++ converge faster. Additionally, DifAttack++ is less prone to overfitting compared to DifAttack, where overfitting leads to a gradual increase in Avg.Q, as depicted in Fig. 3(c). Furthermore, DifAttack++ demonstrates more stable performance on the test set with smaller oscillations, as illustrated in Fig. 3(c) and (d).
@@
4.3 The Open-set Scenario
The experiments above confirms that in closed-set scenarios, DifAttack++ achieves superior ASR and query efficiency compared to compared SOTA methods while maintaining high visual quality of AEs. In this subsection, we will further evaluate attack performance in open-set scenarios, where the victim model’s training set differs from that of surrogate models. We use the challenging task of attacking CLIP models trained on large-scale datasets and a real-world API with unknown training data as examples.
4.3.1 Black-box Attacks against CLIP models.
Take the untargeted attack as an example, we generate AEs for test images from food101 and objectNet datasets to attack their corresponding CLIP models. The surrogate models are Swin-V2-T, ConvNeXtBase, and ResNet101 pretrained on ImageNet. As shown in Table 5, it is worth noting that we only evaluat attack methods applicable for such an open-set scenario. The results demonstrate that DifAttack++ achieves significant improvements in ASR and Avg.Q over other black-box methods, especially when the victim model and classification task are challenging, e.g. classification on ObjectNet via ViT-H-14-CLIPA-336 model. Both the methods based on feature disentanglement, DifAttack and DifAttack++ perform satisfactorily, with DifAttack++ showing further improvements in query efficiency over DifAttack.
Table 6: The ASR and Avg.Q when attacking Imagga API.
@@ 4.3 The Open-set Scenario
4.3.2 Black-box Attacks against Real-world API.
To evaluate the attack performance in a more practical open-set scenario, we assess comparative methods by attacking the real-world Imagga API Imagga (2024). This API utilizes a recognition model trained on a diverse dataset of daily images spanning over 3,000 unknown categories. The Imagga provides multiple predicted classes and corresponding scores for each query image. Due to the API’s call limits, we choose the top-performing six methods in the above experiments as the comparative methods, which are DifAttack, MCGSquare, CGA, GFCS, SimBA, $\mathcal{N}$Attack and SignH. We use the autoencoder and surrogate models pre-trained on ImageNet to extract disentangled features of test images from ImageNet. In untargeted attacks, we aim to remove top 3 predicted classes of clean images. In targeted attacks, we set the target class to be the 2nd-highest scoring class among API’s predicted classes. The experimental results in Table 6 indicate that even when the actual training dataset is completely unknown, our DifAttack++ still exhibits superior ASR and Avg.Q simultaneously than comparative methods both in untargeted and targeted attacks. Notably, GFCS and MCGSquare successfully generate AEs in untargeted attacks but not in targeted attacks, since they are unable to craft AEs when the target class does not occur in the output of surrogate models. It is worth noting that due to the possibility of continuous model updates of real-world APIs over time, we retest the comparative methods, which makes the experimental results inconsistent with those in DifAttack.
Table 7: The ASR of AEs from ImageNet against various SOTA defensive techniques in the untargeted setting.
@@
Figure 4: The visualization of disentangled performance of visual and adversarial features. Images $\tilde{\mathbf{x}}_{1}$ and $\tilde{\mathbf{x}}_{2}$ are the reconstructed version of $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ respectively by our pretrained authoencoder $\mathcal{G}$. The image $\tilde{\mathbf{x}}_{v1a2}$ are reconstructed by $\mathcal{G}$ using the VFs of $\mathbf{x}_{1}$ and the AFs of $\mathbf{x}_{2}$, while $\tilde{\mathbf{x}}_{v2a1}$ are reconstructed by $\mathcal{G}$ using the VFs of $\mathbf{x}_{2}$ and the AFs of $\mathbf{x}_{1}$.
Table 8: The ablation study on the usage of cross domain (CD), the HDF module, and a transferable AE $\hat{\mathbf{x}}_{T}$ as the initial image in the attack stage.
@@
4.4 Black-box Attack against Defensive Methods
It is also important to evaluate the performance of attack methods against defensive strategies. We compare the ASR of comparative methods in untargeted attacks when encountering several SOTA defensive techniques, including Adversarially Training (AT) and pre-processing based methods, such as Bit-depth Reduction (Bit-Red) Xu et al (2018), Neural Representation Purifier (NRP) Naseer et al (2020) and Diffusion Purifier (DiffPure) Nie et al (2022). For attacking AT models pre-trained on ImageNet, we adopt adversarially trained ResNet-50 as the surrogate model and the more powerful classifier Wide-ResNet-50-2 as the victim model Salman et al (2020). For other defensive techniques, we use them to process AEs from ImageNet generated for attacking standardly trained ResNet-18, and then evaluate the ASR of these AEs after being processed. The experimental results shown in Table 7 showcase that Our DifAttack++ achieves the highest average ASR among the current four advanced defensive techniques, surpassing the second-place MCGSquare by nearly 10%. DifAttack++ is also the only method among the comparison methods with an average ASR exceeding 50%. Furthermore, DifAttack++ obtains two first places and two second places in terms of ASR. These results indicate that the AEs generated by DifAttack++ are more effective in resisting various types of defense mechanisms compared to other attack methods, and demonstrate the ability of DifAttack++ to effectively decouple adversarial features from adversarially trained models. Looking back at Fig. 2, the adversarial perturbations generated by MCGSquare consist mainly of colorful blocks, which differ from conventional adversarial noise patterns. Consequently, MCGSquare exhibits significant effectiveness against defense mechanisms that focus solely on perturbations created by gradient-based attacks, such as traditional adversarial training, and techniques like the DiffPure that employ diffusion models for denoising. The aforementioned two defensive techniques primarily operate at the pixel level, whereas methods like NRP that perform denoising in deep feature space show promising results against MCGSquare, reducing its ASR to as low as 10%. In constrast, our DifAttack++ still achieve a highest ASR of 48%. Note that to be consistent with previous works Cheng et al (2019); Feng et al (2022), and due to the poor performance, the evaluation of targeted attacks under defense is temporarily omitted.
@@ 4.5 Analysis on Disentangled Features
To gain a more intuitive understanding of the characteristics of VFs and AFs disentangled by our method, we conduct an experiment to observe the changes on reconstructed images by exchanging the VFs and AFs of a pair of images.
Specifically, we exemplify applying a well pre-trained autoencoder $\mathcal{G}$ to reconstruct and disentangle clean images $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, resulting in their respective reconstructed images ($\tilde{\mathbf{x}}_{1}$ and $\tilde{\mathbf{x}}_{2}$) along with their VFs and AFs. Next, we input $\mathbf{x}_{1}$’s VFs and $\mathbf{x}_{2}$’s AFs into the feature fusion layer $\mathbf{M}$ of $\mathcal{G}$, obtaining the mixed image $\tilde{\mathbf{x}}_{v1a2}$. Similarly, using $\mathbf{x}_{1}$’s AFs and $\mathbf{x}_{2}$’s VFs, we have $\tilde{\mathbf{x}}_{v2a1}$.
As shown in Fig. 4, we randomly select three samples from the test dataset to visualize this process. Notably, the visual quality of the reconstructed images is predominantly influenced by VFs, with minimal impact from AFs. For instance, when the VFs used for reconstructing an image originate from $\mathbf{x}_{1}$, the reconstructed image (i.e., $\tilde{\mathbf{x}}_{1}$ and $\tilde{\mathbf{x}}_{v1a2}$) closely resembles $\mathbf{x}_{1}$ in visual quality no matter where their AFs come from. Similarly, VFs from $\mathbf{x}_{2}$ yield $\tilde{\mathbf{x}}_{2}$ and $\tilde{\mathbf{x}}_{v2a1}$, which closely resemble $\mathbf{x}_{2}$. On the other hand, by observing the residual in reconstructed images when only AFs are replaced while keeping VFs constant (i.e., $\tilde{\mathbf{x}}_{1}-\tilde{\mathbf{x}}_{v1a2}$ and $\tilde{\mathbf{x}}_{2}-\tilde{\mathbf{x}}_{v2a1}$), we find that AFs significantly influence the region of category-specific objects in the reconstructed images, akin to the classifier’s attention. For example, in the first row of samples, when $\tilde{\mathbf{x}}_{v1a2}$’s AFs come from $\mathbf{x}_{2}$, the residual image $\tilde{\mathbf{x}}_{1}-\tilde{\mathbf{x}}_{v1a2}$ focuses on the “black stork” object, which corresponds to the brightest area (large pixel values) in the image $\tilde{\mathbf{x}}_{1}-\tilde{\mathbf{x}}_{v1a2}$. Conversely, when $\tilde{\mathbf{x}}_{v2a1}$’s AFs come from $\mathbf{x}_{1}$, the changes concentrate on the “dining table” object (highlighted in $\tilde{\mathbf{x}}_{2}-\tilde{\mathbf{x}}_{v2a1}$). This phenomenon is also similarly evident in the other two rows of samples.
Based on the experimental results and analysis above, we can draw two conclusions: 1) Our DifAttack++ effectively disentangles the VFs and AFs of images. 2) VFs primarily determine the visual quality of images, and are mainly dependent on images’ signals themselves, while AFs are more associated with the attention regions of the classifier and relatively independent on images’ signals.
@@ 4.6 Ablation Studies
In this subsection, we conduct ablation experiments on three key improvements introduced in our proposed DifAttack++ method compared to our previous work DifAttack: 1) the training stage transitioned from using a single autoencoder for both clean images and AEs to two separate autoencoders for clean and adversarial domains. This Cross Domain is denoted as CD; 2) the decoupling module within the autoencoder changed from a single-layer decouple-fusion (DF) module to a HDF module; and 3) the initial images at the attack stage shifted from clean samples to transferable AEs in close-set scenarios, namely from $\mathbf{x}$ to $\hat{\mathbf{x}}_{T}$ generated by SI-NI-TI-DIM. The evaluation results of these upgrades are summarized in Table 8. To assess the advantages of CD, we maintain the autoencoder architecture used in DifAttack and retaining clean initial images, while adopting two autoencoders responsible for clean and adversarial domains separately, denoted as Dif+CD. Subsequently, to evaluate the HDF, we replace the DF in each autoencoder of Dif+CD with HDF, denoted as Dif+CD+HDF. Finally, to assess the gain of transferable AEs, we replace the clean initial images $\mathbf{x}$ in Dif+CD+HDF with the AE $\hat{\mathbf{x}}_{T}$, representing our ultimate DifAttack++ method used in closed-set scenarios.
Table 8 shows that the three new upgrades proposed in DifAttack++ effectively enhance the ASR and query efficiency when attacking different victim models. Consequently, we can infer the following: 1) distinguishing between clean and adversarial domains helps autoencoders focus more on learning and disentangling image features in the current domain without confusion from the other domain’s distribution; 2) HDF provides more independent disentangled features compared to single-layer DF; 3) utilizing AFs of transferable AEs as a starting point can more efficiently identify successful AFs compared to using AFs from clean samples.

### v3

@@ 4.1 Experiment Setup
Table 4: The visual quality comparison of AEs on ImageNet. AEs are generated by DifAttack++ or MCGSquare.
@@ 4.2 The Close-set Scenario
Tables 2 and 3 present the results on ASR and average query numbers (Avg.Q) when attacking victim models on ImageNet for the “simple group" and “complex group", respectively. For targeted attacks, we randomly set a target class (864) and 1,000 images correctly classified by the victim model, excluding those belonging to the target class, as the test set. It is demonstrated that on average (see the “Average” columns), our DifAttack++ surpasses all compared methods in both ASR and Avg.Q. More specifically, when compared with the second-best method in the challenging targeted attacks, i.e. MCGSquare, for the “simple group" (see Table 2), DifAttack++ reduces Avg.Q by 37.4% meanwhile improving the ASR by an average of 1%. Regarding results for “complex group" in Table 3, DifAttack++ achieves a higher ASR (1.1%) and decreases the Avg.Q from 3,733 to 2,619 (29.8% reduction). Similarly, in untargeted attacks, DifAttack++ diminishes the Avg.Q by 41.5% and 27.5% for the “complex group" and “simple group", respectively, when compared with MCGSquare.
@@
Besides the overall improvement of the attack efficiency and ASR, DifAttack++ also significantly outperforms MCGSquare in terms of the visual quality of AEs, making DifAttack++ less perceptible and more practical. As can be seen from Fig. 2, AEs generated by MCGSquare results in severe visual distortions, e.g., colored blocks or stripes, which are due to the algorithmic characteristics of SquareAttack and MCGSquare. In fact, MCGSquare based on SquareAttack produces similar patterns of adversarial perturbations as SquareAttack does. More severely distorted AEs can be observed in Andriushchenko et al (2020). In contrast, our DifAttack++ crafts AEs with much subtle and imperceptible perturbations. This is because we effectively disentangle the visual and adversarial features of images, and ensure that images’ VFs remain almost unchanged during the generation of AEs, while updating only the AFs. Quantitatively, we also evaluate the similarity between AEs and their associated clean images for DifAttack++ and MCGSquare. From the results summarized in Table 4, we notice that the quality of AEs’ generated by DifAttack++ significantly surpasses that of MCGsquare for all considered metrics; e.g. over 2.5 dB gains in PSNR.
In addition to the above comparisons with SOTA attack methods, we would also like to compare with our early version DifAttack. From Tables 2-3, it can be seen that DifAttack++ outperforms DifAttack by big margins, in terms of ASR and query efficiency. Also, DifAttack++ greatly enhances the training efficiency and test stability. In Fig. 3, we depict the curve of test Avg.Q with respect to the training time, under untargeted attacks against four different victim models. As shown in Figs. 3(a) and (b), it is indicated that, generally, models trained by DifAttack++ converge faster. For example, when the victim model is EfficientNet-B3, using an NVIDIA RTX 3090 GPU and a batch size of 16, DifAttack++ only requires 4 hours to obtain two well-trained autoencoders, whereas DifAttack takes approximately 1 day. Additionally, DifAttack is more likely to overfit to surrogate models, as DifAttack leads to a gradual increase in the test Avg.Q with the prolongation of training time, as depicted in Fig. 3(c). In contrast, DifAttack++ demonstrates more stable attack performance with smaller oscillations, as illustrated in Fig. 3(a)-(d).
Figure 3: The test Avg.Q with respect to the training time, when using DifAttack or DifAttack++ to perform untargeted attacks on ImageNet.
4.3 The Open-set Scenario
The experiments above confirm that in closed-set scenarios, DifAttack++ achieves superior ASR and query efficiency compared to SOTA methods, while still maintaining high visual quality of AEs. In this subsection, we further evaluate the attack performance in the more challenging open-set scenarios, where the victim model’s training set could differ from those of surrogate models. Here, we consider the challenging task of attacking CLIP models trained on large-scale datasets and a real-world API with unknown training data.
4.3.1 Black-box Attacks against CLIP models
Taking the untargeted attack as an example, we generate AEs for test images from Food101 and ObjectNet datasets to attack their corresponding CLIP models. The employed surrogate models are Swin-V2-T, ConvNeXtBase, and ResNet101 pretrained on ImageNet. As shown in Table 5, it is worth noting that we only evaluate attack methods applicable for such an open-set scenario. The results demonstrate that DifAttack++ achieves significant improvements in ASR and Avg.Q over other black-box methods, especially when the victim model and classification task are relatively complicated, e.g. classification on ObjectNet via ViT-H-14-CLIPA-336 model. In fact, both methods based on feature disentanglement, namely, DifAttack and DifAttack++ perform satisfactorily, with DifAttack++ showing further improvements in query efficiency.
@@ 4.3.1 Black-box Attacks against CLIP models
Table 6: The ASR and Avg.Q when attacking real-world Imagga API.
@@
4.3.2 Black-box Attacks against Real-world API
To evaluate the attack performance in a more practical open-set scenario, we assess comparative methods by attacking the real-world Imagga API Imagga (2024). This API utilizes a recognition model trained on a diverse dataset of daily images spanning over 3,000 unknown categories. The Imagga provides multiple predicted classes and corresponding scores for each query image. Due to the API’s call limits, we choose the top-performing six methods in the above experiments as the comparative methods, which are DifAttack, MCGSquare, CGA, GFCS, SimBA, $\mathcal{N}$Attack and SignH. We use the autoencoder and surrogate models pre-trained on ImageNet to extract disentangled features of test images from ImageNet. In untargeted attacks, we aim to remove the top 3 predicted classes of clean images; while in targeted attacks, we set the target class to be the second-highest scoring class among API’s predicted classes. The experimental results in Table 6 suggest that even when the actual training dataset is completely unknown, our DifAttack++ still exhibits superior ASR and Avg.Q simultaneously than comparative methods both in untargeted and targeted attacks. Notably, GFCS and MCGSquare successfully generate AEs in untargeted attacks but not in targeted attacks, since they are unable to craft AEs when the target class does not occur in the output of surrogate models.
Figure 4: The visualization of disentangled representation in DifAttack++. Images $\tilde{\mathbf{x}}_{1}$ and $\tilde{\mathbf{x}}_{2}$ are the reconstructed versions of $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$ respectively by our pretrained authoencoder $\mathcal{G}$. The image $\tilde{\mathbf{x}}_{v1a2}$ is reconstructed by $\mathcal{G}$ using the VFs of $\mathbf{x}_{1}$ and the AFs of $\mathbf{x}_{2}$, while $\tilde{\mathbf{x}}_{v2a1}$ is reconstructed by $\mathcal{G}$ using the VFs of $\mathbf{x}_{2}$ and the AFs of $\mathbf{x}_{1}$.
4.4 Black-box Attack against Defensive Methods
It is also important to evaluate the performance of attack methods against defensive strategies. We compare the ASR of comparative methods in untargeted attacks when encountering several SOTA defensive techniques, including Adversarially Training (AT) and pre-processing based methods, such as Bit-depth Reduction (Bit-Red) Xu et al (2018), Neural Representation Purifier (NRP) Naseer et al (2020) and Diffusion Purifier (DiffPure) Nie et al (2022). For attacking AT models pre-trained on ImageNet, we adopt adversarially trained ResNet-50 as the surrogate model and the more powerful classifier Wide-ResNet-50-2 as the victim model Salman et al (2020). For other defensive techniques, we use them to process AEs from ImageNet generated for attacking standardly trained ResNet-18, and then evaluate the ASR of these AEs after being processed. The experimental results shown in Table 7 showcase that our DifAttack++ achieves the highest average ASR against these four advanced defensive techniques, surpassing the second-place MCGSquare by nearly 10%. DifAttack++ is also the only method, among the comparative ones, with an average ASR exceeding 50%. Furthermore, DifAttack++ obtains two first places and two second places in terms of ASR. These results indicate that the AEs generated by DifAttack++ are more effective in resisting various types of defense mechanisms. Meanwhile, it demonstrates the ability of DifAttack++ to effectively decouple adversarial features from adversarially trained models.
Another interesting phenomenon is from Fig. 2, where we notice that the adversarial perturbations generated by MCGSquare consist mainly of colorful blocks; quite differ from conventional adversarial noise patterns. Consequently, MCGSquare exhibits desirable performance against the defense mechanism that focuses mainly on perturbations created by gradient-based attacks, e.g., the traditional AT, or DiffPure, which is more suitable for dealing with irregular adversarial noise patterns due to its diffusion process.
Table 7: The ASR of AEs from ImageNet against various SOTA defensive techniques in untargeted settings.
@@ 4.4 Black-box Attack against Defensive Methods
Table 8: The ablation study on the usage of CD, the HDF module, the adoption of the initial image, and the number of disentanglement layers.
@@ 4.5 Analysis on Disentangled Features
To gain a more intuitive understanding of the characteristics of VFs and AFs disentangled by our method, we would like to observe the changes in reconstructed images by exchanging the VFs and AFs of a pair of images. Specifically, we first apply a well pre-trained autoencoder $\mathcal{G}$ to reconstruct and disentangle clean images $\mathbf{x}_{1}$ and $\mathbf{x}_{2}$, resulting in their respective reconstructed images ($\tilde{\mathbf{x}}_{1}$ and $\tilde{\mathbf{x}}_{2}$) along with their VFs and AFs. Next, we input $\mathbf{x}_{1}$’s VFs and $\mathbf{x}_{2}$’s AFs into the feature fusion layer $\mathbf{M}$ of $\mathcal{G}$, obtaining the mixed image $\tilde{\mathbf{x}}_{v1a2}$. Similarly, using $\mathbf{x}_{1}$’s AFs and $\mathbf{x}_{2}$’s VFs, we have $\tilde{\mathbf{x}}_{v2a1}$.
As shown in Fig. 4, we randomly select three samples from the test dataset to visualize the above process. Notably, the visual quality of the reconstructed images is predominantly influenced by VFs, with minimal impact from AFs. For instance, when the VFs used for reconstructing an image originate from $\mathbf{x}_{1}$, the reconstructed image (i.e., $\tilde{\mathbf{x}}_{1}$ and $\tilde{\mathbf{x}}_{v1a2}$) closely resembles $\mathbf{x}_{1}$ in visual quality, no matter where their AFs come from. Similarly, VFs from $\mathbf{x}_{2}$ yield $\tilde{\mathbf{x}}_{2}$ and $\tilde{\mathbf{x}}_{v2a1}$, which closely resemble $\mathbf{x}_{2}$. On the other hand, by observing the residual in reconstructed images when only AFs are replaced while keeping VFs unaltered (i.e., $\tilde{\mathbf{x}}_{1}-\tilde{\mathbf{x}}_{v1a2}$ and $\tilde{\mathbf{x}}_{2}-\tilde{\mathbf{x}}_{v2a1}$), we find that AFs significantly impact the region of category-specific objects in the reconstructed images, akin to the classifier’s attention. For example, in the first row, when $\tilde{\mathbf{x}}_{v1a2}$’s AFs come from $\mathbf{x}_{2}$, the residual image $\tilde{\mathbf{x}}_{1}-\tilde{\mathbf{x}}_{v1a2}$ focuses on the “black stork” object, which corresponds to the brightest area (large pixel values) in the image $\tilde{\mathbf{x}}_{1}-\tilde{\mathbf{x}}_{v1a2}$. Conversely, when $\tilde{\mathbf{x}}_{v2a1}$’s AFs come from $\mathbf{x}_{1}$, the changes concentrate on the “dining table” object (highlighted in $\tilde{\mathbf{x}}_{2}-\tilde{\mathbf{x}}_{v2a1}$). Such a phenomenon can also be similarly observed in the other two rows.
Based on the experimental results and analysis above, we can draw two conclusions: 1) Our DifAttack++ is capable of effectively disentangling VFs and AFs of images, and 2) VFs primarily determine the visual quality of images, and are mainly dependent on images’ signals themselves, while AFs are more associated with the attention regions of the classifier and relatively independent of images’ signals.
@@ 4.6 Ablation Studies
We now conduct ablation experiments on the number of disentanglement layers in HDF and three key modules introduced in our proposed DifAttack++ (relative to our early version DifAttack). Specifically, These three modules are 1) CD: the training stage transitioned from using a single autoencoder for both clean images and AEs to two separate autoencoders for clean and adversarial domains, i.e. cross-domain; 2) HDF: the disentanglement module within the autoencoder changed from a single-layer DF module to a two-layer HDF module; and 3) $\hat{\mathbf{x}}_{T}$: the initial images at the attack stage shifted from clean samples to transferable AEs in close-set scenarios, namely from $\mathbf{x}$ to $\hat{\mathbf{x}}_{T}$ generated by SI-NI-TI-DIM. The evaluation results are summarized in Table 8. To assess the advantages of CD, we maintain the autoencoder architecture used in DifAttack and retain clean initial images, while adopting two autoencoders responsible for clean and adversarial domains separately, denoted as Dif+CD. Subsequently, to evaluate the HDF, we replace the DF in each autoencoder of Dif+CD with HDF, denoted as Dif+CD+HDF. To measure the gain of transferable AEs, we replace the clean initial images $\mathbf{x}$ in Dif+CD+HDF with the AE $\hat{\mathbf{x}}_{T}$, representing our ultimate DifAttack++ used in closed-set scenarios. In addition, to further examine the influence of the increasing number of disentanglement layers, we add an extra disentanglement layer behind “Disentanglement layer 2" in Fig. 1(a), while keeping other training and attack schemes the same with Dif+CD+HDF (two layers). The additional layer disentangles input features into AFs and VFs and then fuses these AFs and VFs before feeding them to $\mathbf{M}_{v}$/$\mathbf{M}_{a}$/$\mathbf{M}_{v}^{*}$/$\mathbf{M}_{a}^{*}$. This comparative module is designated as “Dif+CD+HDF(3L)".
Table 8 shows that the three new modules proposed in DifAttack++ effectively enhance the average ASR and query efficiency simultaneously when attacking four different victim models in both targeted and untargeted attacks. By noticing the rows of “DifAttack" and “Dif+CD", we see that incorporating CD leads to a 0.5% increase in the average targeted ASR while reducing the average Avg.Q by 6.6%. Also, upon the introduction of the HDF, i.e. from “Dif+CD" to “Dif+CD+HDF", the average targeted ASR improves by 0.1% while the average Avg.Q decreases by 9.2%. Besides, $\hat{\mathbf{x}}_{T}$ also brings nearly 1,000 (from 2,150 to 1,211) reduction in terms of average targeted Avg.Q. As the untargeted attack is a relatively simpler task compared to the targeted attack and DifAttack already achieves a 100% ASR and an Avg.Q of around 100, the enhancement effects of various modules on untargeted attacks fluctuate slightly across different victim models. But on average, the settings of CD, HDF, and $\hat{\mathbf{x}}_{T}$ still reduce the average Avg.Q by 23/19/34, respectively. Moreover, we observe that more layers of disentanglement do not necessarily result in better performance. A comparison between “Dif+CD+HDF" and “Dif+CD+HDF(3L)" reveals that introducing a third layer not only increases model parameters but also provides limited performance improvement. For instance, although the average ASR in the untargeted and targeted attacks remains stable or slightly increases by 0.1%, the average Avg.Q remains comparable and does not decrease noticeably. This may be because two disentanglement layers are already adequate for the current dataset and classification task, while additional layers could potentially lead to the model overfitting.